# Towards Realistic Evaluation of Commit Message Generation by Matching Online and Offline Settings

### v1

@@ IV-C Results & Discussion
TABLE II: Spearman correlation coefficients $Q(m)$ as defined in Equation 1. The metrics are sorted by the absolute values of correlation coefficients.
@@
The only metric $m$ showing high correlation with the values $ED(G_{i},E_{i})$ that represent the online quality of the CMG system is edit distance. We hypothesize that the reason for edit distance having the highest correlation among the metrics considered lies in the model-generated messages $G$ in our dataset being notably more verbose than both their edited versions $E$ and original messages $O$. In an expert-labeled subset of our dataset, the average number of characters for $G$ is $636$, while for $E$ and $O$, it is $282$ and $204$, respectively.
@@
In a moderately correlated group, we observe only edit similarity. As edit similarity is a normalized version of edit distance, similar reasoning applies: likely, this metric correlates with $ED(G,E)$ due to model-generated messages $G$ being much longer than both $E$ and $O$ in our dataset.
@@
Surprisingly, we observe rather low correlations for widely used metrics for reference-based evaluation in CMG research works, such as BLEU, METEOR, and ROUGE. ChrF and BERTScore, which were not often applied to CMG, are also in this group. Another finding is that, while larger values indicate closer similarity for all the metrics in this group (unlike edit distance, where the smaller values are better), we still observe a positive correlation for ChrF and METEOR in Table II. However, for BLEU, ChrF, and METEOR, our findings are not statistically significant at the confidence level of 95%. Overall, the applicability of metrics from this group remains unclear, as the results are mixed.
Observation 3. BERTScore, ROUGE, BLEU, ChrF and METEOR show low correlation with the online quality of the CMG system and thus might not be the best choice for reference-based CMG evaluation in a setup similar to ours.
@@ V Limitations
Another limitation is that we conducted experiments for a single CMG model, a proprietary LLM GPT-4 in a zero-shot setup. Hence, the findings we obtained may not transfer to CMG systems producing significantly different commit messages. Our key priority was to follow the up-to-date setup from the CMG system in the AI Assistant JetBrains product.
Finally, we note that the expert-labeled dataset we collected is relatively small, with only 57 manually edited samples. To be able to make reliable conclusions, we synthetically extend the dataset via an LLM. While synthetically generated data may not be perfect, it is a common step employed by numerous researchers and practitioners [14]. Additionally, we verify that the distribution of synthetically obtained samples remains close to both the expert-labeled dataset and the PyCharm user logs. Moreover, we describe our pipeline in detail and make all the artifacts publicly available to facilitate further research. [^4] Our HuggingFace collection: https://jb.gg/cmg-evaluation

### v2

@@ IV-C Results & Discussion
TABLE II: Spearman correlation coefficients $Q(m)$ as defined in Equation 1. The metrics are sorted by the absolute values of correlation coefficients. Note that our online metric is edit distance, where larger values indicate greater differences. For all considered metrics $m$ (except edit distance), it is not the case, hence, negative correlation is expected.
@@
The only metric $m$ showing high correlation with the values $ED(G_{i},E_{i})$ that represent the online quality of the CMG system is edit distance. We hypothesize that the reason for edit distance in an offline evaluation setting having the highest correlation with edit distance in an online evaluation setting lies in the model-generated messages $G$ in our dataset being notably more verbose than both their edited versions $E$ and original messages $O$. In an expert-labeled subset of our dataset, the average number of characters for $G$ is $636$, while for $E$ and $O$, it is $282$ and $204$, respectively.
@@
In a moderately correlated group, we observe only edit similarity. As edit similarity is a normalized version of edit distance, similar reasoning applies: likely, this metric correlates with $ED(G_{i},E_{j}),E_{j}\in\text{related}(G_{i})$ due to model-generated messages $G$ being much longer than both $E$ and $O$ in our dataset.
@@
Surprisingly, we observe rather low correlations for widely used metrics for reference-based evaluation in CMG research works, such as BLEU, METEOR, and ROUGE. ChrF and BERTScore, which were not often applied to CMG, are also in this group. Another finding is that, while larger values indicate closer similarity for all the metrics in this group (unlike our primary online metric, edit distance, for which the smaller values are better), we still observe a positive correlation with $ED(G_{i},E_{j}),E_{j}\in\text{related}(G_{i})$ for ChrF and METEOR in Table II. However, for BLEU, ChrF, and METEOR, our findings are not statistically significant at the confidence level of 95%. Overall, the applicability of metrics from this group remains unclear, as the results are mixed.
Observation 3. BERTScore, ROUGE, BLEU, ChrF, and METEOR show low correlation with the online quality of the CMG system and thus might not be the best choice for reference-based CMG evaluation.
@@ V Limitations
Another limitation is that we conducted experiments for a single CMG model, a proprietary LLM GPT-4, in a zero-shot setup. Our key priority in model selection was to follow the up-to-date setup from the CMG system in the JetBrains AI Assistant product. The findings we obtained may not transfer to CMG systems producing significantly different commit messages. For example, models generating less verbose messages might yield different results. Still, GPT-4 is a powerful LLM with strong performance across numerous tasks, including commit message generation [11], making it potentially representative beyond the specific JetBrains feature.
Next, we propose the framework for selecting the most suitable text similarity metric for offline experiments based on available logs before deploying a new version of the model. To further verify the framework, the next step would be to deploy the new version of the model and confirm that the findings from offline experiments remain relevant to the new logs from real users. Deployment to production and conducting online experiments are time-consuming processes and thus fall outside the scope of this work.
Finally, we note that the expert-labeled dataset we collected is relatively small, with only 57 manually edited samples. To be able to make reliable conclusions, we synthetically extend the dataset via an LLM. While synthetically generated data may not be perfect, it is a common step employed by numerous researchers and practitioners [14]. Additionally, we verify that the distribution of synthetically obtained samples remains close to both the expert-labeled dataset and the PyCharm user logs. Moreover, we describe our pipeline in detail and make all the artifacts publicly available to facilitate further research.